# Heat current and fluctuations between a dissipative qubit and a monitor under continuous measurement and feedback

### v2

@@ 5 Summary
We study the statistical property of heat exchange between the monitor and the dissipative qubit under continuous measurement and feedback. The first half is the steady-state heat current when changing the measurement and feedback states. The heat current can take a wider range of values compared with the measurement-only case. When doing the measurement to the ground state and the feedback to the excited state, heat maximally flows from the monitor to the qubit, which exceeds the maximum value for the measurement-only case. We also observe the qubit cooling by the measurement and the feedback, which is never seen for the measurement-only case. We identify the boundary between the cooling and the heating is determined by the temperature, the cooling region is extended as the temperature decreases.
The second part of this paper is devoted to the fluctuation around the heat current by unraveling the Lindblad equation, the stochastic master equation. The power spectrum is composed of the frequency-independent and -dependent contributions, coming from the quantum jump and the measurement backaction dynamics, called the Poisson noise and the backaction noise, respectively, in this work. Both noises are characteristic of the continuous measurement process, different from the standard shot noise such as in the electronic circuit. For a large range of parameters, these spectra have the opposite sign, which can be interpreted as the measurement backaction and leads to the sub-Poisson Fano factor. However, around zero heat change by the quantum jump, these spectra have the same sign due to the dissipation to the heat baths, resulting in the super-Poisson Fano factor.

### v3

@@ 5 Summary
We study the statistical property of energy exchange between the monitor and the dissipative qubit under continuous measurement and feedback. The first half is the steady-state energy flow when changing the measurement and feedback states. The energy flow can take a wider range of values compared with the measurement-only case. When doing the measurement to the ground state and the feedback to the excited state, energy maximally flows from the monitor to the qubit, which exceeds the maximum value for the measurement-only case. We also observe the qubit cooling by the measurement and the feedback, which is never seen for the measurement-only case. We identify the boundary between the cooling and the heating is determined by the temperature, the cooling region is extended as the temperature decreases.
The second part of this paper is devoted to the fluctuation of the energy flow by unraveling the Lindblad equation, the stochastic master equation. The power spectrum is composed of the frequency-independent and -dependent contributions, coming from the quantum jump and the measurement backaction dynamics, called the Poisson noise and the backaction noise, respectively, in this work. Both noises are characteristic of the continuous measurement process, different from the standard shot noise such as in the electronic circuit. For a large range of parameters, these spectra have the opposite sign, which can be interpreted as the measurement backaction and leads to the sub-Poisson Fano factor. However, around zero energy change by the quantum jump, these spectra have the same sign due to the dissipation to the heat baths, resulting in the super-Poisson Fano factor.